# From Source to Target: Leveraging Transfer Learning for Predictive Process Monitoring in Organizations

### v1

@@ 5.1 Initial model building on source
The first phase, which aims to build and evaluate an initial base model $f_{\text{DNN}}(\cdot)$ in the source context $S$, consists of the following six steps.
@@
Second, the activities and the timestamps of the events of the event log $L_{S}$ are encoded. The activities are encoded by transforming their textual values into high-dimensional feature vector representations using a pre-trained embedding model $f_{\text{EMB}}:\sum^{*}\rightarrow\mathrm{R}^{n}$, where $\sum$ is the alphabet, $n$ is the embedding size, and $\sum^{*}$ are all finite activities that can be built from the alphabet $\sum$. For example, common pre-trained models are word2vec (Mikolov et al. 2013), pre-trained global vectors for word representation (GloVe) (Pennington et al. 2014), and BERT (Devlin et al. 2019) models. In doing so, activity names are represented more abstractly than, for example, in the case of a one-hot encoding, facilitating the transfer from a source to a target domain – that is, symmetrical, feature-based TL. Depending on the architecture and capability of the pre-trained embedding model used, the feature vectors representing the values of the activity attribute have different sizes. For example, a word2vec model is a small neural network model with one hidden layer, which learns the relationship between words and produces feature vectors of sizes, such as 100 and 300.121212https://radimrehurek.com/gensim/models/word2vec.html (Accessed 23 July 2025) In contrast, a BERT model is a DNN with several hidden layers (e.g., 12 layers), that learns the meaning of a word in a sentence depending on the surrounding words (context) and produces larger feature vectors, for example, of size 768.131313https://github.com/google-research/bert (Accessed 23 July 2025)
@@
Fourth, the prefixes of the event log $L_{S}$ are transformed into the event data $\mathsf{X}_{S}\in\mathrm{R}^{s\times T\times v}$ and the label data $\textbf{y}_{S}\in\mathrm{R}^{s}$ to build a DNN-based base model. Here, $s$ is the number of prefixes, $T$ refers to the number of time steps of the longest prefix, and $v$ denotes the number of features. For the event data $\mathsf{X}_{S}$, the remaining space for prefixes comprising fewer time steps than the prefix with the maximum number of time steps is padded with zeros. For each time step of each prefix, $\mathsf{X}_{S}$ stores information on the activity and the timestamp attribute in the form of extracted features. The concatenation of the activity features and the time features represents the final feature vector with its size corresponding to the number of features $v$ of the event data $\mathsf{X}_{S}$. For the label data, labels are set on the trace level to predict target values as early as possible. This is common in outcome-oriented PPM (Teinemaa et al. 2019), and applicable in our work, as we focus on binary, outcome-based predictions. For example, if the prediction target of interest is the target in-time, the model is applied at each time step of a running process instance to predict whether this instance will be finished in time or not. Finally, the event data $\mathsf{X}_{S}$ and the label data $\mathbf{y}_{S}$ are split into a training set $(\mathsf{X}_{S}^{tr},\mathbf{y}_{S}^{tr})$ and a validation set $(\mathsf{X}_{S}^{val},\mathbf{y}_{S}^{val})$ for model training, and a test set $(\mathsf{X}_{S}^{te},\mathbf{y}_{S}^{te})$ for model evaluation.
Fifth, given the training set of the source domain $(\mathsf{X}_{S}^{tr},\mathbf{y}_{S}^{tr})$, the base model $f_{\text{DNN}}:\mathsf{X}_{i}\in\mathrm{R}^{T\times v}\rightarrow\hat{\mathbf{y}_{i}}\in(0,1)$ is created and trained, mapping event data instances $\mathsf{X}_{i}=(\mathbf{x}_{1},\dots,\mathbf{x}_{T})$, with $x_{t}\in\mathrm{R}^{v}$ onto predictions $\hat{\mathbf{y}}_{i}$. The base model is a two-layer LSTM model. For each layer $l$ = {1,2} and time step $t=1,\dots,T$, the LSTM (Hochreiter and Schmidhuber 1997) transition is
@@
$(\mathbf{h}_{T}^{(2)},\mathbf{c}_{T}^{(2)})=f_{\text{LSTM}}(\mathsf{X_{i}}),$ (7)
@@
Then, the hidden state $\mathbf{h}_{T}^{(2)}$ is taken, projected through a fully-connected layer, and a sigmoid activation is applied to the projection’s output to calculate the prediction, as formalized in Eq. (8).
$\mathbf{\hat{y}}=sigmoid(W_{\text{out }}\mathbf{h}_{T}^{(2)}+\mathbf{b}_{\text{out }}).$ (8)
@@
Before base model optimization, the model’s internal parameters are initialized to improve the learning from the event data of the source context and to facilitate knowledge transfer to the target context. All input weight matrices $\mathbf{W}_{\bigoplus}$ for $\bigoplus\in\{f,i,g,o\}$ are Xavier-uniform initialized to center the values around 0, going into sigmoid/tanh activations. This enables the LSTM to start in a flexible area and encourages the optimizer to obtain useful gradients immediately. All recurrent weight matrices $\mathbf{U}_{\bigoplus}$ for $\bigoplus\in\{f,i,g,o\}$ are initialized with a random orthogonal matrix to keep the singular values of the recurrent map near 1 and to reduce the problem of exploiting/vanishing gradients (Arjovsky et al. 2016). In the recurrent part of the LSTM, the recurrent map is the operation that takes the hidden state at t1t-1 and transforms it to the next hidden state at time $t$, using the recurrent weight matrices. Singular values of the recurrent weight matrices measure how much the recurrent map stretches or squashes vectors in different directions. For example, if a singular value is >1>1, that direction grows exponentially over time (i.e., gradients explode). The forget-gate bias $\mathbf{b}_{f}$ is set to +1 (other LSTM biases $\{\mathbf{b}_{i},\mathbf{b}_{g},\mathbf{b}_{o}\}$ are set to 0) to encourage the network to preserve memory at the start of training (Jozefowicz et al. 2015).
For the final layer, the weights $\mathbf{W}_{out}$ are He-uniform initialized with 0 bias to allow for more stability in training, given the fact that our model has one output neuron to model the two possible process-related outcomes (e.g., in-time or not), which our approach assumes.
@@
The internal parameters of the base model for prediction are optimized by solving the following optimization problem:
$\beta^{*}_{S}$ $=\underset{\beta_{S}}{\arg\min}\sum^{s}_{i=1}\pazocal{L}\Big{(}f_{{\text{DNN}}}\Big{(}\mathsf{X}_{S,i}^{tr};\beta_{S}\Big{)},\mathbf{y}_{S,i}^{tr}\Big{)},$ (9)
@@
where $\pazocal{L}$ is a binary cross entropy loss (BCE loss) function, $\beta^{*}_{S}$ are the adjusted weights of the model after training on training event data of the source context $S$, $s$ is the number of prefixes, $\mathsf{X}_{S,i}^{tr}$ and $\mathbf{y}_{S,i}^{tr}$ are the i-th prefix and label of the training event data $\mathsf{X}^{tr}_{S}$ and the training label data $\mathbf{y}_{S}^{tr}$ of the source context $S$. During model optimization, early stopping is performed if no improvement in the BCE loss on the validation event data $(\mathsf{X}_{S}^{val},\mathbf{y}_{S}^{val})$ takes place over 10 subsequent epochs.
@@
Sixth, the built base model $f_{\text{DNN}}(\cdot)$ is applied to the test set $(\mathsf{X}_{S}^{te},\mathbf{y}_{S}^{te})$ and its prediction performance is measured via ML metrics $m\in M$; that is, $m(f_{\text{DNN}}(\mathsf{X}_{S}^{te};\beta^{*}_{S}),\mathbf{y}_{S}^{te}))$. As ML metrics, commonly used metrics in the context of outcome-oriented PPM are calculated, such as the $\text{AUC}_{\text{ROC}}$ and the F1-score (Teinemaa et al. 2019).
@@ 5.2 Transfer resources from source to target
The second phase, which aims to transfer resources from the source context $S$ that are relevant for the target context $T$, consists of two steps.
@@
First, relevant resources are transferred from the source context $S$ to the target context $T$. These resources can be grouped into TL-specific resources and general resources. TL-specific resources include the (pre-trained) embedding model for activity encoding, information for timestamp encoding (one or more selected time features for transfer and divisors for feature value scaling), and the pre-trained prediction model for application. General resources include further information on pre-processing (e.g., the type of prefix encoding) and evaluation (e.g., ML metrics for assessing the model’s prediction performance).
@@
Second, if traces in the form of an event log $L_{T}$ are available of the target context $T$, the base model $f_{\text{DNN}}$ from the source context $S$ can be applied and evaluated in an intermediary step before it is used to create predictions based on (ongoing) process instances in the target context $T$. In this intermediary step, first, the event log $L_{T}$ is loaded, and activity and timestamp encoding are performed. For activity encoding, the transferred, pre-trained embedding model $f_{\text{EMB}}(\cdot)$ is applied, whereas for timestamp encoding, the transferred information on time feature selection and scaling is used. After activity and timestamp encoding, prefixes are encoded, the event data $\mathsf{X}_{T}$ and the label data $\mathbf{y}_{T}$ are created, and the data are split into a training set $(\mathsf{X}_{T}^{tr},\mathbf{y}_{T}^{tr})$, a validation set $(\mathsf{X}_{T}^{val},\mathbf{y}_{T}^{val})$, and a test set $(\mathsf{X}_{T}^{te},\mathbf{y}_{T}^{te})$. To ensure that the pre-processing is performed in the same way as in the source context $S$, further pre-processing information is transferred from the source context $S$ to the target context $T$. Moreover, the transferred model $f_{\text{DNN}(\cdot)}$ could be fine-tuned with event data of the target context $T$ (i.e., $(\mathsf{X}_{T}^{tr},\mathbf{y}_{T}^{tr})$ and $(\mathsf{X}_{T}^{val},\mathbf{y}_{T}^{val})$). However, by default, the base model is not fine-tuned with any event data of the target context. Finally, the transferred base model $f_{\text{DNN}}(\cdot)$ is applied to the test set $(\mathsf{X}_{T}^{te},\mathbf{y}_{T}^{te})$ to evaluate its prediction performance using ML metrics $m\in\pazocal{M}$; that is, $m(f_{\text{DNN}}(\mathsf{X}_{T}^{te};\beta^{*}_{S}),\mathbf{y}_{T}^{te})$. For this purpose, evaluation information, such as ML metrics, is transferred from the source to the target context to assess the prediction performance of the base model in the same way as in the source context $S$.
@@
However, for the evaluation in this intermediary step, there are two further aspects to consider. First, as the test sets of the event logs from the target and source contexts are naturally different, the test ML metrics cannot be directly compared with each other. To overcome that, a second prediction model should be trained “from scratch” using the training set $(\mathsf{X}_{T}^{tr},\mathbf{y}_{T}^{tr})$ of the target context $T$. Based on the test set $(\mathsf{X}_{T}^{te},\mathbf{y}_{T}^{te})$, this second prediction model is compared with the transferred base model in terms of prediction performance to evaluate the performance of the transfer. Second, if the event log $L_{T}$ only comprises a few traces, the entire event log should be used as a test set in the evaluation of the base model to get an initial indication of how well the base model performs in the target context.
@@ 5.3 Online application of model to target
The third phase, which aims to create predictions for a new, ongoing process instance in the target context $T$ based on the resources transferred from the source context $S$, consists of six steps. First, the (ongoing) process instance $\sigma_{T}$ is loaded, which represents a prefix of a trace that will be completed in the future. Then, activity encoding and timestamp encoding are performed on the (ongoing) process instance. For the activity encoding, the transferred, pre-trained embedding model $f_{\text{EMB}}(\cdot)$ is applied, and for timestamp encoding, the transferred information on time feature selection and scaling is used. In the next step, the prefix is encoded before it is transformed into event data $\mathsf{X}_{T}$. To also ensure that the pre-processing in the online application is performed in the same way as in the source context $S$, further pre-processing information is transferred from the source context $S$ to the target context $T$. Then, the transferred base model $f_{\text{DNN}}(\cdot)$ is applied to the event data $\mathsf{X}_{T}$ to produce the prediction $\hat{y}$. In contrast to the previous two phases, the prediction is not evaluated, as no label information is available in the online application. Instead, in this phase, the prediction is used for operational support by process users or other stakeholders to, for example, proactively draw attention to undesired process outcomes, with the overall objective of improving the performance of the underlying business process.
@@ 6 Evaluation
To evaluate our technique, we instantiated it in two real-life use cases. The first one represents an intra-organizational case, whereas the second one is an inter-organizational case. In doing so, we account for the differences between transfer scenarios and can thoroughly evaluate the performance of our technique for both intra- and inter-organizational scenarios. For example, for an intra-organizational use case, to extend the application of an existing prediction model to a new product or service in the same process context, performed activities can be similar, while variations in other dimensions, like sequence of activities or throughput time, can occur. For an inter-organizational use case, the performed activities might even vary in their descriptions.

### v2

@@ 5.1 Initial model building on source
The first phase, which aims to build and evaluate an initial base model $f_{\text{DNN}}(\cdot)$ in the source context $S$, consists of six steps.
@@
Second, the activities and the timestamps of the events of the event log $L_{S}$ are encoded. The activities are encoded by transforming their textual values into high-dimensional feature vector representations using a pre-trained embedding model $f_{\text{EMB}}:\sum^{*}\rightarrow\mathrm{R}^{n}$, where $\sum$ is the alphabet, $n$ is the embedding size, and $\sum^{*}$ are all finite activities that can be built from the alphabet $\sum$. For example, word2vec (Mikolov et al. 2013) and BERT (Devlin et al. 2019) are common natural-language-processing methods to create pre-trained embedding models. By encoding activity names in this way, they are represented more abstractly than, for example, in the case of a one-hot encoding, facilitating the transfer from a source to a target domain – that is, symmetrical, feature-based TL (Zhuang et al. 2020). Depending on the architecture and capability of the pre-trained embedding model used, the feature vectors representing the values of the activity attribute have different sizes. For example, a word2vec model is a small neural network model with one hidden layer, which learns the relationship between words and produces feature vectors of lengths, such as 100 and 300.121212https://radimrehurek.com/gensim/models/word2vec.html (Accessed 23 July 2025) In contrast, a BERT model is a DNN with several hidden layers (e.g., 12 layers), that learns the meaning of a word in a sentence depending on the surrounding words (context) and produces larger feature vectors, for example, of length 768.131313https://github.com/google-research/bert (Accessed 23 July 2025)
@@
Fourth, the prefixes of the event log $L_{S}$ are transformed into the event data $\mathsf{X}_{S}\in\mathrm{R}^{s\times T\times v}$ and the label data $\textbf{y}_{S}\in\mathrm{R}^{s}$ to build a DNN-based base model. Here, $s$ is the number of prefixes, $T$ refers to the number of time steps of the longest prefix, and $v$ denotes the number of features. For the event data $\mathsf{X}_{S}$, the remaining space for prefixes comprising fewer time steps than the prefix with the maximum number of time steps is padded with zeros. For each time step of each prefix, $\mathsf{X}_{S}$ stores information on the activity and the timestamp attribute in the form of extracted features. The concatenation of the activity features and the time features represents the final feature vector with its size corresponding to the number of features $v$ of the event data $\mathsf{X}_{S}$. For the label data, labels are set on the trace level to predict target values as early as possible. This is common in outcome-oriented PPM (Teinemaa et al. 2019), and applicable in our work, as we focus on binary, outcome-based process predictions. For example, if the prediction target of interest is in-time, the model is applied at each time step of a running process instance to predict whether this instance will be finished in time or not. Finally, the event data $\mathsf{X}_{S}$ and the label data $\mathbf{y}_{S}$ are split into a training set $(\mathsf{X}_{S}^{tr},\mathbf{y}_{S}^{tr})$ and a validation set $(\mathsf{X}_{S}^{val},\mathbf{y}_{S}^{val})$ for model training, and a test set $(\mathsf{X}_{S}^{te},\mathbf{y}_{S}^{te})$ for model evaluation.
Fifth, given the training event data of the source context $\mathsf{X}_{S}^{tr}=(\mathbf{X}_{1},\dots,\mathbf{X}_{s})$, with $\mathbf{X}_{i}\in\mathrm{R}^{T\times v}$, the base model $f_{\text{DNN}}:\mathbf{X}_{i}\rightarrow\hat{y}_{i}\in(0,1)$ is created and trained, mapping event data instances $\mathbf{X}_{i}=(\mathbf{x}_{1},\dots,\mathbf{x}_{T})$, with $\textbf{x}_{t}\in\mathrm{R}^{v}$, onto estimations $\hat{y}_{i}$. The base model is a two-layer LSTM model. For each layer $l$ = {1,2} and time step $t=1,\dots,T$, the LSTM (Hochreiter and Schmidhuber 1997) transition is
@@
$\Big(\mathbf{h}_{T}^{(2)},\mathbf{c}_{T}^{(2)}\Big)=f_{\text{LSTM}}\Big(\mathbf{X}_{i}\Big),$ (7)
@@
Then, the hidden state $\mathbf{h}_{T}^{(2)}$ is taken, projected through a fully-connected layer, and a sigmoid activation is applied to the projection’s output to calculate the estimations, as formalized in Eq. (8).
$\hat{y}=sigmoid\Big(W_{\text{out }}\mathbf{h}_{T}^{(2)}+\mathbf{b}_{\text{out }}\Big).$ (8)
@@
Before optimizing the base model’s internal parameters, they are initialized to improve learning from the event data of the source context and to facilitate knowledge transfer to the target context. All input weight matrices $\mathbf{W}_{\bigoplus}$ for $\bigoplus\in\{f,i,g,o\}$ are Xavier-uniform initialized to center the values around 0, going into sigmoid/tanh activations. This enables the LSTM to start in a flexible area and encourages the optimizer to obtain useful gradients immediately. All recurrent weight matrices $\mathbf{U}_{\bigoplus}$ for $\bigoplus\in\{f,i,g,o\}$ are initialized with a random orthogonal matrix to keep the singular values of the recurrent map near 1 and to reduce the problem of exploiting/vanishing gradients (Arjovsky et al. 2016). In the recurrent part of the LSTM, the recurrent map is the operation that takes the hidden state at t1t-1 and transforms it to the next hidden state at time $t$, using the recurrent weight matrices. Singular values of the recurrent weight matrices measure how much the recurrent map stretches or squashes vectors in different directions. For example, if a singular value is >1>1, that direction grows exponentially over time (i.e., gradients explode). The forget-gate bias $\mathbf{b}_{f}$ is set to +1 (other LSTM biases $\{\mathbf{b}_{i},\mathbf{b}_{g},\mathbf{b}_{o}\}$ are set to 0) to encourage the network to preserve memory at the start of training (Jozefowicz et al. 2015). For the final layer, the weights $\mathbf{W}_{out}$ are He-uniform initialized with 0 bias to allow for more stability in training, given the fact that our model has one output neuron to model the two possible process-related outcomes (e.g., in-time or not), which our technique assumes.
@@
The internal parameters of the base model are optimized by solving the following optimization problem:
$\beta^{*}_{S}$ $=\underset{\beta_{S}}{\arg\min}\sum^{s}_{i=1}\pazocal{L}\Big(f_{{\text{DNN}}}\Big(\mathbf{X}_{S,i}^{tr};\beta_{S}\Big),y_{S,i}^{tr}\Big),$ (9)
@@
where $\pazocal{L}$ is a binary cross entropy loss (BCE loss) function, $\beta^{*}_{S}$ are the adjusted internal parameters of the model after training on training event data of the source context $S$, $s$ is the number of prefixes, $\mathbf{X}_{S,i}^{tr}$ is the i-th prefix of the training event data $\mathsf{X}^{tr}_{S}$ of the source context $S$, and $y_{S,i}^{tr}$ is the i-th label value of training label data $\mathbf{y}_{S}^{tr}$ of the source context $S$. During the internal parameter optimization, early stopping is performed if no improvement in the BCE loss on the validation set $(\mathsf{X}_{S}^{val},\mathbf{y}_{S}^{val})$ takes place over 10 subsequent epochs.
@@
Sixth, the built base model $f_{\text{DNN}}(\cdot)$ is evaluated to determine whether it should be transferred from the source context $S$ to the target context $T$. To do that, the base model is applied to the test event data $(\mathsf{X}_{S}^{te},\mathbf{y}_{S}^{te})$ and its prediction performance is measured via ML metrics $m\in M$; that is, $m(u(f_{\text{DNN}}(\mathsf{X}_{S}^{te};\beta^{*}_{S})),\mathbf{y}_{S}^{te})$, where $u$ is a function transforming model estimations to class predictions or probabilities. As ML metrics, commonly used in outcome-oriented PPM are calculated, such as the $\text{AUC}_{\text{ROC}}$ and the F1-score (Teinemaa et al. 2019). In addition, the prediction performance, measured via the ML metrics, should be considered depending on the used validation strategy, as it determines how the event data is split for model evaluation.
@@ 5.2 Transfer resources from source to target
The second phase, which aims to transfer relevant and usable resources from the source context $S$ to the target context $T$, consists of two steps.
@@
First, relevant resources are transferred from the source context to the target context. These resources can be grouped into TL-specific resources and general resources. TL-specific resources include the pre-trained embedding model for activity encoding, information for timestamp encoding (one or more selected time features for transfer and divisors for feature value scaling), and the pre-trained prediction model for application. General resources include further information on pre-processing (e.g., the type of prefix encoding) and evaluation (e.g., ML metrics for measuring the base model’s prediction performance).
@@
Second, if traces in the form of an event log $L_{T}$ are available of the target context $T$, the base model $f_{\text{DNN}}(\cdot)$ from the source context $S$ can be applied and evaluated in an intermediary step before it is used to create class predictions and probabilities based on (ongoing) process instances in the target context $T$. In this intermediary step, first, the event log $L_{T}$ is loaded, and activity and timestamp encoding are performed. For activity encoding, the transferred, pre-trained embedding model $f_{\text{EMB}}(\cdot)$ is applied, whereas for timestamp encoding, the transferred information on time feature selection and scaling is used. After activity and timestamp encoding, prefixes are encoded, the event data $\mathsf{X}_{T}$ and the label data $\mathbf{y}_{T}$ are created, and the data are split into a training set $(\mathsf{X}_{T}^{tr},\mathbf{y}_{T}^{tr})$, a validation set $(\mathsf{X}_{T}^{val},\mathbf{y}_{T}^{val})$, and a test set $(\mathsf{X}_{T}^{te},\mathbf{y}_{T}^{te})$. To ensure that the pre-processing is performed in the same way as in the source context $S$, further pre-processing information is transferred from the source context $S$ to the target context $T$. Moreover, the transferred model $f_{\text{DNN}}(\cdot)$ could be fine-tuned with event data of the target context $T$ (i.e., $(\mathsf{X}_{T}^{tr},\mathbf{y}_{T}^{tr})$ and $(\mathsf{X}_{T}^{val},\mathbf{y}_{T}^{val})$). However, by default, the base model is not fine-tuned with any event data of the target context. Finally, the transferred base model is applied to the test set $(\mathsf{X}_{T}^{te},\mathbf{y}_{T}^{te})$ to evaluate its prediction performance. For this purpose, evaluation information, such as used ML metrics, is transferred from the source context $S$ to the target context $T$ to assess the prediction performance of the base model in the same way as in the source context.
@@
For the evaluation in this intermediary step, there are two further aspects to consider. First, as the test sets of the event logs from the target and source contexts are naturally different, the test ML metrics cannot be directly compared with each other. To overcome that, a second prediction model should be trained “from scratch” using the training set $(\mathsf{X}_{T}^{tr},\mathbf{y}_{T}^{tr})$ of the target context $T$. Based on the test set $(\mathsf{X}_{T}^{te},\mathbf{y}_{T}^{te})$, this second prediction model is compared with the transferred base model in terms of prediction performance to evaluate the performance of the transfer. Second, if the event log $L_{T}$ only comprises a few traces, the entire event log should be used as a test set in the evaluation of the base model to get an initial indication of how well the base model performs in the target context.
@@ 5.3 Online application of model to target
The third phase, which aims to create class predictions and probabilities for a new, ongoing process instance in the target context $T$ based on the resources transferred from the source context $S$, consists of six steps.
First, the (ongoing) process instance $\sigma_{T}$ is loaded, which represents a prefix of a trace that will be completed in the future. Then, activity encoding and timestamp encoding are performed on the (ongoing) process instance. For the activity encoding, the transferred, pre-trained embedding model $f_{\text{EMB}}(\cdot)$ is applied, and for timestamp encoding, the transferred information on time feature selection and scaling is used. In the next step, the prefix is encoded before it is transformed into the event data instance $\mathbf{X}_{T}$. To also ensure that the pre-processing in the online application is performed in the same way as in the source context $S$, further pre-processing information is transferred from the source context $S$ to the target context $T$. Then, the transferred base model $f_{\text{DNN}}(\cdot)$ is applied to the event data instance $\mathbf{X}_{T}$ to produce the estimate $\hat{y}$, which is transformed into a class prediction and distribution. In contrast to the previous two phases, the prediction performance is not evaluated, as no label information is available in the online application. Instead, in this phase, the class prediction and distribution are used for operational support by process users or other stakeholders to, for example, proactively draw attention to undesired process outcomes, with the overall objective of improving the performance of the underlying business process.
@@ 6 Evaluation
To evaluate our technique, we instantiated it in two real-life use cases. The first one represents an intra-organizational use case, whereas the second one is an inter-organizational use case. In doing so, we account for the differences between transfer scenarios and can thoroughly evaluate the performance of our technique for both intra- and inter-organizational scenarios. For example, in an intra-organizational use case where the application of an existing prediction model is extended to a new product or service within the same process context, the performed activities can be similar, while variations may occur in other dimensions, such as the sequence of activities or throughput time. In an inter-organizational use case, the performed activities might even vary in their descriptions.